# Dynamic Pre-training: Towards Efficient and Scalable All-in-One Image Restoration

### v1

@@ 3.2 Dynamic Pre-training Strategy
In recent times, large-scale pre-training has become a key strategy to improve given network performance. Initializing the network with pre-trained weights provides a robust foundation, boosting performance even with fewer parameters. However, this strategy is resource-intensive, demanding considerable computational power and GPU hours. Therefore, we propose a dynamic pre-training that is capable of simultaneous training of multiple variants of the network. These networks are unified by shared network modules weight but varying network depth. This strategy enables the concurrent training of a diverse set of models tailored to various computational needs and task complexities, all while leveraging a shared baseline architecture. Through the proposed dynamic pre-training strategy, we train both DyNet-L and DyNet-S versions of our DyNet concurrently in a single training session, achieving a 50% reduction in GPU hours.
DyNet-L and DyNet-S utilize the same underlying weights for the transformer blocks initialized at each encoder-decoder level. They just differ in the reuse frequency of these initialized transformer blocks at each encoder-decoder level. Thus, during training iterations, we randomly alternate between DyNet-L and DyNet-S ensuring the optimization of the shared underlying weights as shown in Fig. 3. Moreover, we improve the generalization capability of our DyNet variants by adopting an input masking strategy akin to that proposed in masked auto-encoders [19]. We randomly mask portions of an image and train the DyNet variants to reconstruct these masked regions in a self-supervised fashion. The dataset used for this training is described below.
@@ 4 Million-IRD: A Dataset for Image Restoration
Large-scale pre-training for image restoration effectively demands large-scale, high-quality and high-resolution datasets. The current image restoration datasets [27, 2, 35], when combined, offer only a few thousand images. This is considerably inadequate compared to the extensively large-scale datasets [42, 41] utilized in pre-training for other high-level tasks, such as visual recognition, object detection, and segmentation. The relatively small size of the existing training sets for image restoration restricts the performance capabilities of the underlying networks. More importantly, we are motivated by the scaling laws in high-level tasks that demonstrate the large-scale pre-taining can enable even lightweight, efficient model designs to reach the performance mark of much heavier models [1]. To address this gap, we introduce a new million-scale dataset named Million-IRD having $\sim$2M high-quality, high-resolution images, curated specifically for the pre-training of models for image restoration tasks. Examples from our Million-IRD dataset are shown in the Fig. 4. Our data collection and pre-processing pipeline are discussed below.
@@ 4.2 Data post-processing
Overall, our Million-IRD dataset has 2.09 million high-quality, high-resolution images. Each image in our dataset comes with metadata detailing its download link and the image resolution. A detailed breakdown of images and their respective sources is provided in the supplementary material. We extract high-resolution non-overlapping patches (of spatial size $512^{2}$) from each image, followed by the application of a flat region detector which eliminates any patch comprising more than 50% flat area. This post-processing phase allows us to assemble a pool of $\sim$8 million diverse image patches, tailored for the pre-training phase of our DyNet variants.
@@ 5 Experiments
We validate the proposed DyNet across three key image restoration tasks: dehazing, deraining, and denoising. In line with the existing PromptIR [36], our experiments are conducted under two distinct setups: (a) All-in-One, wherein a single model is trained to handle all three degradation types, and (b) Single-task, where individual models are trained for each specific image restoration task.
@@ 5.1 Implementation Details
Datasets. For all-in-one and single-task implementations, we follow the same training protocol as existing PromptIR [36] and finetune the pre-trained DyNet-S and DyNet-L. Following [26, 36], we prepare datasets for various restoration tasks. Specifically, for image denoising, we combine the BSD400 [5] (400 training images) and WED [29] (4,744 images) datasets, adding Gaussian noise at levels $\sigma\in{15,25,50}$ to create noisy images. Testing is conducted on the BSD68 [30] and Urban100 [20] datasets. For deraining, we use Rain100L [50] with 200 training and 100 testing clean-rainy image pairs. Dehazing employs the SOTS [25] dataset, featuring 72,135 training and 500 testing images. To develop a unified model for all tasks, we merge these datasets, fine-tune pre-trained DyNet-L and DyNet-S for 120 epochs and directly evaluate them across different tasks. While, for a single task, the pre-trained DyNet-L is fine-tunned for 120 epochs on a respective training set of the task.
Table 1: Comparison results in the All-in-one restoration setting. Our DyNet-L model outperforms PromptIR [36] by 0.82 dB on average across tasks. Additionally, our DyNet-S model achieves a 0.59 dB average improvement over PromptIR [36], with reductions of 31.34% in parameters and 56.75% in GFlops.
Comparative Dehazing Deraining Denoising on BSD68 dataset [30]) Average
Methods on SOTS [25] on Rain100L [16] $\sigma=15$ $\sigma=25$ $\sigma=50$
BRDNet [44] 23.23/0.895 27.42/0.895 32.26/0.898 29.76/0.836 26.34/0.836 27.80/0.843
LPNet [17] 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738
FDGAN [12] 24.71/0.924 29.89/0.933 30.25/0.910 28.81/0.868 26.43/0.776 28.02/0.883
MPRNet [54] 25.28/0.954 33.57/0.954 33.54/0.927 30.89/0.880 27.56/0.779 30.17/0.899
DL[16] 26.92/0.391 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875
AirNet [26] 27.94/0.962 34.90/0.967 33.92/0.933 31.26/0.888 28.00/0.797 31.20/0.910
PromptIR [36] 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913
DyNet-S (Ours) 31.51/0.980 38.11/0.981 34.06/0.935 31.41/0.891 28.15/0.802 32.65/0.918
DyNet-L (Ours) 31.98/0.981 38.71/0.983 34.11/0.936 31.44/0.892 28.18/0.803 32.88/0.920
Figure 5: Comparative analysis of image denoising by all-in-one methods on the BSD68 dataset [30]. Our DyNet-L reduces noise, producing more sharp and clear image compared to the PromptIR [36].
@@ 5.2 Comparisons on multiple degradations under All-in-one setting
We benchmark our DyNet against a range of general image restoration models [44, 17, 12, 54] and specific all-in-one solutions [16, 26, 36], as shown in Table 1. On average, across various tasks, DyNet-L and DyNet-S outperform the previously best PromptIR [36] by $0.82$ dB and $0.59$ dB, respectively, with DyNet-S also cutting down on 56.75% of parameters and 31.34% of GFlops. For image denoising, DyNet-S achieves a $1.09$ dB higher average PSNR compared to DL[16], with Fig. 5 showcasing its ability to deliver noise-free images with improved structural integrity. Notably, DyNet-L sets a new benchmark in image deraining and dehazing with improvements of $2.34$ dB and $1.4$ dB in PSNR, respectively. Visual comparisons in Fig. 6 and  7 demonstrate DyNet’s superior capability in removing rain and haze, respectively, producing more cleaner images than PromptIR [36].
Figure 6: Comparative analysis of image deraining by all-in-one methods on the Rain100L dataset [16]. Our DyNet-L successfully eliminates rain streaks, producing clear, rain-free images as compared to the recent PromptIR [36].
Figure 7: Comparative analysis of image dehazing by all-in-one methods on the SOTS dataset [25]. Our approach reduces haze, producing more clear image compared to the PromptIR [36].
Table 2: Dehazing results in the single-task setting on the SOTS benchmark dataset [25]. Our DyNet-L achieves a boost of $0.76$ dB over PromptIR [26].
DehazeNet[6] MSCNN[39] AODNet[24] EPDN[37] FDGAN[12] AirNet[26] Restormer[52] PromptIR [36] DyNet-L
22.46/0.851 22.06/0.908 20.29/0.877 22.57/0.863 23.15/0.921 23.18/0.900 30.87/0.969 31.31/0.973 32.07/0.982
@@ 5.3 Comparisons on single degradation
We assess DyNet-L’s efficacy in single-task settings, where distinct models are tailored to specific restoration tasks, demonstrating the effectiveness of content-adaptive prompting through prompt blocks. Our results, as shown in Table 2, indicate DyNet-L’s superior performance with an improvement of $0.76$ dB over PromptIR [36] and $1.2$ dB over Restormer [52] in dehazing. This pattern is consistent across other tasks, including deraining and denoising. Notably, compared to PromptIR [36], DyNet-L achieves performance gains of $1.81$ dB in deraining (refer to Table 3) and achieves a $0.13$ dB improvement in denoising at a noise level of $\sigma=50$ on the Urban100 dataset (Table 4).
Table 3: Deraining results in the single-task setting on Rain100L [16]. Compared to the PromptIR [36], our method yields $1.68$ dB PSNR improvement.
DIDMDN[55] UMR[51] SIRR[48] MSPFN[21] LPNet[17] AirNet[26] Restormer[52] PromptIR DyNet-L
@@
Table 4: Denoising comparisons in the single-task setting on BSD68 [30] and Urban100 [20] datasets. For the challenging noise level of $\sigma=50$ on Urban100 [20], our DyNet-L obtains $0.64$ dB gain compared to AirNet [26].
Comparative BSD68 [30] Urban100 [20]
Methods $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$
CBM3D [10] 33.50/0.922 30.69/0.868 27.36/0.763 33.93/0.941 31.36/0.909 27.93/0.840
DnCNN [57] 33.89/0.930 31.23/0.883 27.92/0.789 32.98/0.931 30.81/0.902 27.59/0.833
IRCNN [58] 33.87/0.929 31.18/0.882 27.88/0.790 27.59/0.833 31.20/0.909 27.70/0.840
FFDNet [59] 33.87/0.929 31.21/0.882 27.96/0.789 33.83/0.942 31.40/0.912 28.05/0.848
BRDNet [44] 34.10/0.929 31.43/0.885 28.16/0.794 34.42/0.946 31.99/0.919 28.56/0.858
AirNet [26] 34.14/0.936 31.48/0.893 28.23/0.806 34.40/0.949 32.10/0.924 28.88/0.871
Restormer [52] 34.29/0.937 31.64/0.895 28.41/0.810 34.67/0.969 32.41/0.927 29.31/0.878
PromptIR [36] 34.34/0.938 31.71/0.897 28.49/0.813 34.77/0.952 32.49/0.929 29.39/0.881
DyNet-L (Ours) 34.40/0.938 31.76/0.897 28.53/0.813 34.84/0.952 32.60/0.930 29.52/0.882
@@
Comparative Dehazing Deraining Denoising on BSD68 dataset [30]) Average Par GFlops
Methods on SOTS  [25] on Rain100L [16] $\sigma=15$ $\sigma=25$ $\sigma=50$
PromptIR [36] 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913 37M 242.355
@@
DyNet-S 31.51/0.980 38.11/0.981 34.06/0.935 31.41/0.891 28.15/0.802 32.65/0.918 16M 166.38
DyNet-L 31.98/0.981 38.71/0.983 34.11/0.936 31.44/0.892 28.18/0.803 32.88/0.920 16M 242.35
@@ 5.4 Ablation studies
To study the impact of different components, we conduct various ablation experiments in the all-in-one setting, as summarized in Table 5. Table 5(a) illustrates that our DyNet-L and DyNet-S achieves comparable performance to the baseline PromptIR [36] with reductions of 56.75% in parameters and 31.34% in GFlops, without the proposed dynamic pre-training on Million-IRD dataset. This shows that the proposed fundamental correction of placing prompt block at skip connections is effective compared to on the decoder side as given in PromptIR [36]. Furthermore, pre-training on our Million-IRD dataset brings a significant improvement, resulting in an average PSNR increase of 0.55 dB for DyNet-L compared to the DyNet-L without pre-training as shown in Table 5(b). A similar effect is observed for our DyNet-S variant. Moreover, our proposed dynamic training strategy reduces GPU hours by 50% when training variants of the DyNet. Overall, each of our contributions (i.e., implicit prompting via skip connections, dynamic network architecture, Million-IRD dataset, and dynamic pre-training strategy) synergistically enhances the performance of the proposed approach, while significantly reducing GFlops and learning parameters compared to the baseline PromptIR [36].
@@
Table 6: Performance of the proposed DyNet-S on different combinations of degradation types i.e., removal of noise, rain, and haze.
Degradation Denoising on BSD68 dataset [30] Deraining on Dehazing on
Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100L [16] SOTS [25]
34.30/0.938 31.65/0.897 28.42/0.813 - -
- - - 38.72/0.984 -
- - - - 31.45/0.980
34.19/0.936 31.54/0.894 28.30/0.808 38.61/0.984 -
34.11/0.935 31.45/0.893 28.18/0.802 - 31.69/0.981
- - - 38.16/0.982 31.55/0.980
34.06/0.935 31.41/0.891 28.15/0.802 38.11/0.981 31.51/0.980

### v2

@@ 3.2 Dynamic Pre-training Strategy
In recent times, large-scale pre-training has become a key strategy to improve given network performance. Initializing the network with pre-trained weights provides a robust foundation, boosting performance even with fewer parameters. However, this strategy is resource-intensive, demanding considerable computational power and GPU hours. Therefore, we propose a dynamic pre-training strategy that enables simultaneous training of multiple network variants with shared module weights but varying depths. This allows for concurrent training of diverse models tailored to different computational needs and task complexities, all leveraging a shared architecture. Through the proposed dynamic pre-training strategy, we train both DyNet-L and DyNet-S variants of our DyNet concurrently in a single training session, achieving a 50% reduction in GPU hours.
DyNet-L and DyNet-S utilize the same underlying weights for transformer blocks, differing only in the reuse frequency of these blocks at each encoder-decoder level. Thus, during training iterations, we randomly alternate between DyNet-L and DyNet-S ensuring the optimization of the shared underlying weights as shown in Fig. 3. Moreover, we improve the generalization capability of our DyNet variants by adopting an input masking strategy akin to that proposed in masked auto-encoders (He et al., 2022). We randomly mask portions of an image and train the DyNet variants to reconstruct masked regions in a self-supervised fashion. The training dataset is described below.
@@ 4 Million-IRD: A Dataset for Image Restoration
Large-scale pre-training for image restoration effectively demands large-scale, high-quality and high-resolution datasets. The current image restoration datasets: LSDIR (Li et al., 2023a), NTIRE (Agustsson & Timofte, 2017), Flickr2K (Online, ), when combined, offer only a few thousand images. This is considerably inadequate compared to the extensively large-scale datasets LAION-5B (Schuhmann et al., 2022), ImageNet-21K (Ridnik et al., 2021) utilized in pre-training for other high-level tasks, such as visual recognition, object detection, and segmentation. The relatively small size of the existing training sets for image restoration restricts the performance capabilities of the underlying networks. More importantly, we are motivated by the scaling laws in high-level tasks that demonstrate the large-scale pre-taining can enable even lightweight, efficient model designs to reach the performance mark of much heavier models (Abnar et al., 2021). To address this gap, we introduce a new million-scale dataset named Million-IRD having $\sim$2M high-quality, high-resolution images, curated specifically for the pre-training of models for IR tasks. Fig. 4 shows few samples from our Million-IRD dataset. Our data collection and pre-processing pipeline are discussed below.
@@ 4.2 Data post-processing
Overall, our Million-IRD dataset has 2.09 million high-quality, high-resolution images. Each image in our dataset comes with metadata detailing its download link and the image resolution. A detailed breakdown of images and their respective sources are provided in the Appendix A. We extract high-resolution non-overlapping patches (of spatial size $512^{2}$) from each image, followed by the application of a flat region detector which eliminates any patch comprising more than 50% flat area. This post-processing phase allows us to assemble a pool of $\sim$8 million diverse image patches, tailored for the pre-training phase of our DyNet variants.
@@ 5 Experiments
We validate the proposed DyNet across three key image restoration tasks: dehazing, deraining, and denoising. In line with the existing PromptIR (Potlapalli et al., 2023), our experiments are conducted under two distinct setups: (a) All-in-One, wherein a single model is trained to handle all three degradation types, and (b) Single-task, where individual models are trained for each specific image restoration task.
@@ 5.1 Implementation Details
Datasets. For the all-in-one task implementations, we adopt the training protocol from PromptIR and concurrently fine-tune the pre-trained DyNet-S and DyNet-L in a single session. Following PromptIR, we prepare datasets for various restoration tasks. Specifically, for image denoising, we combine the BSD400 (Arbelaez et al., 2011) (400 training images) and WED (Ma et al., 2016) (4,744 images) datasets, adding Gaussian noise at levels $\sigma\in[15,25,50]$ to create noisy images. Testing is conducted on the BSD68 (Martin et al., 2001) and Urban100 (Huang et al., 2015) datasets. For deraining, we use Rain100L (Yang et al., 2020) with 200 training and 100 testing clean-rainy image pairs. Dehazing employs the SOTS (Li et al., 2018) dataset, featuring 72,135 training and 500 testing images. To build a unified model for all tasks, we merge the datasets and deploy a dynamic finetunning of the pre-trained DyNet for 120 epochs, then evaluate it across multiple tasks. The resulting flexible DyNet allows ”on-the-go” adjustments to model depth, switching between DyNet-L and DyNet-S without needing separate model weights. Both versions share the same underlying weights, differing only in depth. For individual tasks, the pre-trained DyNet-L is fine-tuned for 120 epochs on the corresponding task-specific training set.
Table 1: Comparison results in the All-in-one restoration setting. Our DyNet-L model outperforms PromptIR by 0.68 dB on average across tasks. Further, our DyNet-S model achieves a 0.43 dB average improvement over PromptIR, with reductions of 31.34% in parameters and 56.75% in GFlops.
Comparative Dehazing Deraining Denoising on BSD68 dataset Average
Methods on SOTS on Rain100L $\sigma=15$ $\sigma=25$ $\sigma=50$ PSNR/SSIM
BRDNet (Tian et al., 2020) 23.23/0.895 27.42/0.895 32.26/0.898 29.76/0.836 26.34/0.836 27.80/0.843
LPNet (Gao et al., 2019) 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738
FDGAN (Dong et al., 2020b) 24.71/0.924 29.89/0.933 30.25/0.910 28.81/0.868 26.43/0.776 28.02/0.883
MPRNet (Zamir et al., 2021) 25.28/0.954 33.57/0.954 33.54/0.927 30.89/0.880 27.56/0.779 30.17/0.899
DL (Fan et al., 2019) 26.92/0.391 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875
AirNet (Li et al., 2022) 27.94/0.962 34.90/0.967 33.92/0.933 31.26/0.888 28.00/0.797 31.20/0.910
PromptIR (Potlapalli et al., 2023) 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913
DyNet-S (Ours) 30.80/0.980 38.02/0.982 34.07/0.935 31.41/0.892 28.15/0.802 32.49/0.918
DyNet-L (Ours) 31.34/0.980 38.69/0.983 34.09/0.935 31.43/0.892 28.18/0.803 32.74/0.920
Figure 5: Comparative analysis of image denoising by all-in-one methods on the BSD68 and Urban100 dataset. DyNet reduces noise, produces a sharp and clear images compared to the PromptIR
Figure 6: Comparative analysis within all-in-one methods for image de-raining on the Rain100L dataset. Our DyNet-L outperforms PromptIR by producing clearer, rain-free images.
Figure 7: Comparative analysis of image dehazing by all-in-one methods on the real-world hazy images. Our approach reduces haze, producing more clear image compared to the PromptIR.
@@ 5.2 Comparisons on multiple degradations under All-in-one setting
We benchmark our DyNet against a range of general IR methods and specific all-in-one solutions, as shown in Table 1. On average, across various tasks, DyNet-L and DyNet-S outperform the previously best PromptIR by $0.68$ dB and $0.43$ dB, respectively, and DyNet-S also cuts down 56.75% of parameters and 31.34% of GFlops. For image denoising, DyNet-L achieves a $1.11$ dB higher average PSNR compared to DL, and Fig. 5 showcasing its ability to deliver noise-free images with improved structural integrity. Notably, DyNet-L sets a new benchmark in image deraining and dehazing with improvements of $2.32$ dB and $0.76$ dB in PSNR, respectively. In addition, the visual comparisons in Fig. 6 demonstrate DyNet’s superior capability to remove rain, producing cleaner images than PromptIR. For image dehazing, existing methods such as PromptIR are limited to removing neutral color (gray) haze. However, in real-world conditions, haze can appear in different colors, such as light blue during the morning or yellow in smog-filled environments, etc. As illustrated in Fig.  7, PromptIR struggles to restore images with nongray haze due to its rigidity to haze-color types. Considering different types of haze (increasing the prompt length) during training could potentially address this issue. However, considering all types of haze color is not a practical solution. Therefore, to enhance our DyNet’s ability to restore different haze-color images, we first restore the color balance using the existing Gray-World Assumption (GWA) algorithm followed by the dehazing through our network, which effectively recovers the haze-free image. We present a visual comparison between PromptIR and our DyNet on a variety of real-world hazy images in Fig. 7. For a fair assessment, we also include results from GWA+PromptIR (PromptIR+). The results clearly demonstrate that our DyNet outperforms PromptIR in reducing haze effects. The simple plug-in GWA module enables our approach to dehaze real-world hazy images without increasing model complexity.
@@ 5.3 Comparisons on single degradation
We assess DyNet-L’s efficacy in single-task settings, where distinct models are tailored to specific restoration tasks, demonstrating the effectiveness of content-adaptive prompting through prompt blocks. Our results, as shown in Table 2, indicate DyNet-L’s superior performance with an improvement of $0.76$ dB over PromptIR and $1.2$ dB over Restormer in dehazing. This pattern is consistent across other tasks, including deraining and denoising. Notably, compared to PromptIR, DyNet-L achieves performance gains of $1.81$ dB in deraining (refer to Table 3) and achieves a $0.24$ dB improvement in denoising at a noise level of $\sigma=50$ on the Urban100 dataset (Table 4).
Table 2: Dehazing results in the single-task setting on the SOTS (Li et al., 2018) dataset. Our DyNet-L achieves a boost of $0.76$ dB over PromptIR.
MSCNN AODNet EPDN FDGAN AirNet Restormer PromptIR DyNet-L
(Ren et al., 2020) (Cai et al., 2016) (Qu et al., 2019) (Dong et al., 2020b) (Li et al., 2022) (Zamir et al., 2022) (Potlapalli et al., 2023) (Ours)
22.06/0.908 20.29/0.877 22.57/0.863 23.15/0.921 23.18/0.900 30.87/0.969 31.31/0.973 32.07/0.982
Table 3: Deraining results in the single-task setting on Rain100L (Yang et al., 2020). Compared to the PromptIR, our method yields 1.81 dB PSNR improvement.
DIDMDN UMR SIRR MSPFN LPNet AirNet Restormer PromptIR DyNet-L
(Zhang & Patel, 2018) (Yasarla & Patel, 2019) (Wei et al., 2019) (Jiang et al., 2020) (Gao et al., 2019) (Li et al., 2022) (Zamir et al., 2022) (Potlapalli et al., 2023) (Ours)
@@
Table 4: Denoising comparisons in the single-task setting on BSD68 and Urban100 datasets. For difficult noise level of $\sigma=50$ on Urban100, our DyNet-L obtains $0.75$ dB gain compared to AirNet.
Comparative Urban100( (Huang et al., 2015)) BSD68( (Martin et al., 2001)) Average
Methods $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ PSNR/SSIM
CBM3D (Dabov et al., 2007) 33.93/0.941 31.36/0.909 27.93/0.840 33.50/0.922 30.69/0.868 27.36/0.763 30.79/0.874
DnCNN (Zhang et al., 2017a) 32.98/0.931 30.81/0.902 27.59/0.833 33.89/0.930 31.23/0.883 27.92/0.789 30.74/0.878
IRCNN (Zhang et al., 2017b) 27.59/0.833 31.20/0.909 27.70/0.840 33.87/0.929 31.18/0.882 27.88/0.790 29.90/0.864
FFDNet (Zhang et al., 2018) 33.83/0.942 31.40/0.912 28.05/0.848 33.87/0.929 31.21/0.882 27.96/0.789 31.05/0.884
BRDNet (Tian et al., 2020) 34.42/0.946 31.99/0.919 28.56/0.858 34.10/0.929 31.43/0.885 28.16/0.794 31.44/0.889
AirNet (Li et al., 2022) 34.40/0.949 32.10/0.924 28.88/0.871 34.14/0.936 31.48/0.893 28.23/0.806 31.54/0.897
Restormer (Zamir et al., 2022) 34.67/0.969 32.41/0.927 29.31/0.878 34.29/0.937 31.64/0.895 28.41/0.810 31.79/0.903
PromptIR (Potlapalli et al., 2023) 34.77/0.952 32.49/0.929 29.39/0.881 34.34/0.938 31.71/0.897 28.49/0.813 31.87/0.902
DyNet-L (Ours) 34.91/0.953 32.68/0.930 29.63/0.884 34.34/0.938 31.71/0.896 28.47/0.812 31.96/0.902
@@
Comparative Dehazing Deraining Denoising on BSD68 dataset (Martin et al., 2001) Average Par GFlops
Methods on SOTS on Rain100L $\sigma=15$ $\sigma=25$ $\sigma=50$
PromptIR 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913 37M 242.355
@@
DyNet-S 30.80/0.980 38.02/0.982 34.07/0.935 31.41/0.892 28.15/0.802 32.49/0.918 16M 166.38
DyNet-L 31.34/0.981 38.69/0.983 34.09/0.935 31.43/0.892 28.18/0.803 32.74/0.920 16M 242.35
Table 6: Performance of the proposed DyNet-S on different combinations of degradation types.
Degradation Denoising on BSD68 dataset Deraining on Dehazing on
Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100L SOTS
34.28/0.938 31.64/0.897 28.42/0.813 - -
- - - 38.69/0.983 -
- - - - 30.72/0.980
34.18/0.936 31.52/0.893 28.31/0.808 38.54/0.983 -
34.11/0.935 31.44/0.893 28.18/0.802 - 31.08/0.981
- - - 38.12/0.982 30.92/0.980
34.07/0.935 31.41/0.892 28.15/0.802 38.02/0.982 30.80/0.980
@@ 5.4 Ablation studies
To study the impact of different components, we conduct various ablation experiments in the all-in-one setting, as summarized in Table 5. Table 5(a) illustrates that our DyNet-L and DyNet-S achieves comparable performance to the baseline PromptIR with reductions of 56.75% in parameters and 31.34% in GFlops, without the proposed dynamic pre-training on Million-IRD dataset. This shows that the proposed fundamental correction of placing prompt block at skip connections is effective compared to on the decoder side as given in PromptIR. Furthermore, pre-training on our Million-IRD dataset brings a significant improvement, resulting in an average PSNR increase of 0.41 dB for DyNet-L compared to the DyNet-L without pre-training as shown in Table 5(b). A similar effect is observed for our DyNet-S variant. Moreover, our proposed dynamic training strategy reduces GPU hours by 50% when training variants of the DyNet. Overall, each of our contributions (i.e., implicit prompting via skip connections, dynamic network architecture, Million-IRD dataset, and dynamic pre-training strategy) synergistically enhances the performance of the proposed approach, while significantly reducing GFlops and learning parameters compared to the baseline PromptIR.